### v6

# Physics-Aware Semi-Supervised Underwater Image Enhancement

Abstract
Underwater images normally suffer from degradation due to the transmission medium of water bodies. Both traditional prior-based approaches and deep learning-based methods have been used to address this problem. However, the inflexible assumption of the former often impairs their effectiveness in handling diverse underwater scenes, while the generalization of the latter to unseen images is usually weakened by insufficient data. In this study, we leverage both the physics-based underwater Image Formation Model (IFM) and deep learning techniques for Underwater Image Enhancement (UIE). To this end, we propose a novel Physics-Aware Dual-Stream Underwater Image Enhancement Network, i.e., PA-UIENet, which comprises a Transmission Estimation Steam (T-Stream) and an Ambient Light Estimation Stream (A-Stream). This network fulfills the UIE task by explicitly estimating the degradation parameters of the IFM. We also adopt an IFM-inspired semi-supervised learning framework, which exploits both the labeled and unlabeled images, to address the issue of insufficient data. Our method performs better than, or at least comparably to, eight baselines across five testing sets in the degradation estimation and UIE tasks. This should be due to the fact that it not only can model the degradation but also can learn the characteristics of diverse underwater scenes.
1 Introduction
The images captured in the underwater environment play important roles in ocean exploration. However, these images normally suffer from different degradation, due to the wavelength-dependent light absorption and light scattering caused by the underwater transmission medium. According to [1], the underwater Image Formation Model (IFM) can be expressed as:
$I^{c}(x)=J^{c}(x)t^{c}(x)+(1-t^{c}(x))A^{c},$ (1)
where $I^{c}(x)$ ($c\in\{R,G,B\}$) is a degraded image, $J^{c}(x)$ is the underlying clean image, $t^{c}(x)=exp(-\beta^{c}d(x))$ denotes the transmission maps, $\beta^{c}$ is the attenuation factor for each channel, $d(x)$ is the scene distance map and $A^{c}$ is the ambient light. (See Fig.1 for examples). In this context, $J^{c}(x)t^{c}(x)$ describes the distance-dependent color distortion, while $(1-t^{c}(x))A^{c}$ represents the scattering of the ambient light which decreases the contrast and lessens the visibility of the image.
Figure 1: Examples of a degraded underwater image, the corresponding ambient light image $A^{c}$ and three transmission maps $t^{c}(x)$ ($c\in\{R,G,B\}$) estimated using our method, and the enhanced image produced using the estimated parameters on top of the IFM.
Considering the degradation may interfere with downstream tasks, Underwater Image Enhancement (UIE) is key to underwater image processing and machine vision. Due to the complicated degradation process, this task is challenging. Many traditional methods [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] have been proposed, to address the challenge. These methods can be treated as the prior-based methods because they typically rely on a certain form of prior information. Some prior-based methods [2, 3, 4] aimed to estimate the parameters of the IFM. However, they normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene violates the prior assumption [4]. On the other hand, the other prior-based methods [9, 10, 11] directly enhance the quality of underwater images without explicitly modeling the degradation process. Due to the lack of the knowledge of underwater scenes, they cannot effectively enhance the images captured in the complex scenes and often introduce artifacts.
In contrast, deep learning-based UIE approaches [14, 15, 16, 17] aim to directly learn from different types of degraded images. Therefore, the problem with the limited static prior can be alleviated. Ideally, a large number of real-world labeled training images should be utilized. However, it is difficult to collect such a large labeled underwater data set in practice. Alternatively, the best enhanced image can be selected as the reference (ground-truth) by observers from the images processed by a set of UIE algorithms [16, 18]. Nevertheless, this sort of data sets cannot represent diverse underwater scenes due to the limited size. In this case, the deep UIE model trained may overfit the training data.
The synthetic images generated using the IFM [19, 14] or a Generative Adversarial Network (GAN) [20, 21, 22] were employed in order to address the problem with the lack of labeled data. However, the utilization of these images to simulate real-world underwater degradation processes is struggling. Transfer learning techniques [21, 15] were also applied for that purpose, which usually used generative models to map degraded images to the clearer images. But they are difficult to train and may produce unstable results. Recently, promising results were derived using zero-shot learning approaches [23, 24] by directly learning degradation parameters from unlabeled degraded images.
To effectively model the degradation process, we propose a Physics-Aware Dual-Stream Underwater Image Enhancement Network, i.e., PA-UIENet, by injecting the physical principle into the deep neural network. This network performs the UIE task by explicitly estimating the degradation parameters of the IFM using a Transmission Estimation Steam (T-Stream) and an Ambient Light Estimation Stream (A-Stream). We further adopt an IFM-inspired semi-supervised learning framework, including a bi-directional supervised scheme and an unsupervised scheme, to overcome the limitation of insufficient data. Compared with the unsupervised methods [23, 24], the framework can be more easily trained due to the facilitation from the supervised scheme. Since the unsupervised scheme uses additional unlabeled images while the bi-directional supervised scheme exploits labeled images more effectively, it can be better trained than the supervised approaches [14, 15, 16, 17]. As a result, the generalization of the model trained using our framework is stronger than that trained using a supervised or unsupervised method. To our knowledge, this study makes the first attempt to jointly apply the physics-aware deep network and the IFM-inspired semi-supervised learning technique to the UIE task.
Our main contributions can be summarized as threefold. (1) We propose a physics-aware dual-stream network for the UIE task by explicitly estimating the degradation parameters of the IFM. (2) An IFM-inspired semi-supervised learning framework is adopted, to overcome the challenge of insufficient data. (3) We conduct a series of comparative experiments on five real-world underwater testing sets along with eight baselines. These results provide the community with a set of benchmarks.
2 Related Work
2.1 Prior-Based Methods
To recover clear underwater images, many methods [2, 3, 4, 5, 6, 7, 8] used the prior information to estimate the degradation parameters of the IFM. Although the Dark Channel Prior (DCP) [2] was initially proposed for image dehazing, some researchers [5, 6] employed it for underwater image enhancement, due to the similarity between the degradation processes resulted from the foggy weather and underwater environment. However, the original DCP usually yielded erroneous estimations. Therefore, more studies were performed to improve it, such as Underwater Dark Channel Prior (UDCP) [3] and Generalization of the Dark Channel Prior (GDCP) [4]. Moreover, the Haze Line Prior (HL) [7] was utilized in some studies [8, 25]. Despite these methods were designed on top of the IFM, they normally made a rigid assumption about the underwater environment, which restricted the application of these to a specific underwater scenario.
On the other hand, some approaches used the more general priors [26, 9, 10, 11, 12] to enhance the quality of underwater images without taking the IFM into account, including Histogram Equalization (HE) [26, 9], Retinex-based methods [11, 12] and image fusion techniques [13, 27]. These approaches usually enhanced the contrast and produced more color-balanced results. However, they probably struggle with processing globally inhomogeneous degraded images and may introduce artifacts, such as halos and color casts, due to the lack of the knowledge of underwater scenes.
2.2 Learning-Based Methods
Thanks to the powerful representation learning ability and large-scale training data [28], deep learning based methods greatly boosted the development of computer vision in both high-level [29, 30, 31] and low-level vision tasks [32, 33, 34]. However, the application of deep learning to underwater image enhancement is much less than other tasks. This dilemma should be attributed to the difficulty on collecting a large number of underwater images and the corresponding clean counterparts. Some deep learning methods [19, 35] used the fake underwater images, synthesized based on terrestrial images, to train a network for degradation parameter estimation. Since the synthesis algorithms were relatively simple, they usually could not simulate diverse underwater scenes. Besides, the content of the terrestrial images was significantly different from underwater scenes. As a result, those methods usually encountered the domain-shift problem when they were applied to real underwater images.
To overcome this problem, some labeled real-world underwater image data sets, e.g., UIEB [16] and SUIM-E [18], were collected. Regarding a degraded image, a set of enhancement methods were applied. The most visually pleasant result was manually picked out as the reference of the degraded image. A UIE model can be trained using these data sets without considering the IFM. For example, the WaterNet [16] used CNNs to learn fusion weights for the purpose of fusing the enhanced results produced by the White Balance [13], HE and Gamma Correction techniques. To perform the UIE task, the UColor [17] method combined the RGB, HSV and Lab color spaces with the guidance of the IFM parameter estimated using the GDCP [4]. Since the training data sets were relatively small, these methods usually overfitted the training data and were poorly generalized to unseen images. In addition, they could not explicitly exploit the information provided by the IFM.
In contrast to the lack of labeled underwater images, it is easier to collect unlabeled underwater images. Existing studies [15, 36] sourced real-world underwater images and categorized them according to the degree of degradation. Unsupervised learning approaches were developed on top of these data by learning the mapping from severely degraded images to slightly degraded images. However, these approaches were hard to train and tended to produce unstable results. Recently, a semi-supervised UIE method [37] was developed. Nevertheless, this method did not consider physical principles.
The above-mentioned studies either are not robust to diverse underwater scenes, or lack the knowledge of underwater images. We are hence inspired to exploit both the IFM-inspired semi-supervised learning technique and the physics-aware deep neural network to address these issues.
3 Methodology
Considering the importance of physical principles to the UIE task, we propose a Physics-Aware Dual-Stream Underwater Image Enhancement Network, namely, PA-UIENet. This network contains a Transmission Estimation Steam (T-Stream) and an Ambient Light Estimation Stream (A-Stream), which are used to explicitly estimate the parameters of the IFM. To address the challenge of the lack of labeled real-world underwater images, we further adopt an IFM-inspired semi-supervised learning framework, which consists of a bi-directional supervised scheme and an unsupervised scheme. Compared to the supervised or unsupervised method, the PA-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model trained is thus stronger, due to the complementary action of the two schemes.
Figure 2: The architecture of the proposed PA-UIENet, which comprises two individual streams: A-Stream and T-Stream, to estimate the degradation parameters of the IFM.
3.1 PA-UIENet
As illustrated in Fig.2, the PA-UIENet contains an encoder-decoder style T-stream and a Transformer-based A-stream, used to estimate transmission maps and the ambient light respectively. Since the red channel usually encounters the more severe attenuation than the blue and green channels [1, 38], the utilization of this channel is key to color restoration. Therefore, a simple Red Channel Tuner (RCT) module is designed to adaptively emphasize the red channel before an image is fed into the T-stream. The output of the first encoder block Enc_1Enc\_1 in the T-stream is processed by Patchify and the result is sent to the A-stream. A Residual Communication Module (RCM) connects the two blocks of the encoder of the T-stream and the A-stream at the same level, to fulfill the bi-directional feature exchange. Both the transmission maps and the ambient light, estimated using the T-stream and A-stream respectively, are fed into the IFM. As a result, the enhanced underwater image is produced.
Red Chanel Tuner (RCT). Inspired by the channel attention mechanism [39], we propose a compact RCT module in order to adaptively emphasize the information contained in the red channel. This module first uses a convolutional layer to extract basic features from the input image, and then use the Global Average Pooling (GAP) to obtain a global representation. A fully-connected layer and the Sigmoid activation function are further used to transform this representation to a tuning weight. The weight is finally used to scale the red channel of the image.
T-Stream. According to Eq. (1), the transmission map varies at different pixel locations. In this case, the network should be able to generate fine-grained representations. Hence, we build the T-stream using a CNN-based encoder-decoder network. The encoder comprises five blocks, denoted as Enc_1Enc\_1 to Enc_5Enc\_5, and the Max Pooling operation is applied to each block. As a result, multi-scale feature maps are derived using the encoder. Given Enc_5Enc\_5 is used as the bottleneck, the decoder contains four blocks, denoted as Dec_1Dec\_1 to Dec_4Dec\_4, symmetrically. Bilinear interpolation is used to upsample the feature maps within each block. Both the encoder and decoder blocks share the same structure, which consists of two ConvConv-BNBN-ReLUReLU operations. Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure. The output of the last decoder block is fed into a convolutional layer. A set of transmission maps are produced in terms of different color channels.
A-Stream. The ambient light describes the luminance of the scene. Since this parameter does not encode the information of a specific object, it can be treated as homogeneous. We are motivated to design the A-stream using a set of Transformer [40] blocks for the sake of leveraging the self-attention mechanism to extract the global information. Five blocks, denoted as Trans_1Trans\_1 to Trans_5Trans\_5, are comprised of this stream. Each block contains a self-attention layer and a two-layer Multi-layer Perceptron (MLP) unit, in which two residual connections are used respectively. Since the input of the Transformer is a sequence of visual tokens , the feature maps of Enc_1Enc\_1 are processed using Patchify which includes the Flatten and Linear Projection operations. Considering the estimation of the ambient light does not require local characteristics and the quadratic complexity of self-attention, we downsample these maps by a factor of 16 to reduce the computational demand. The result is a set of tokens. A special token, namely, the Ambient token, is also adopted, to specifically learn the ambient light parameter. All the tokens are fed into Trans_1Trans\_1. The Ambient token produced by the last block Trans_5Trans\_5 is sent to a linear layer. Finally, the estimation of the ambient light is achieved.
Residual Communication Module (RCM). According to the IFM, the degradation can be considered as the outcome caused by both the transmission and the ambient light. In this situation, the T-stream and A-stream will learn some common features. Therefore, the feature exchange between both the streams is likely to boost the training of each stream. Recently, the complementary action of CNNs and Transformers has been shown [41]. We are inspired to propose the RCM for the sake of bridging the two streams. Specifically, the tokens produced by a block in the A-stream are folded into a set of feature maps while the feature maps generated by a block of the encoder of the T-stream are downsampled to the shape of these maps. Then both sets of maps are concatenated along the channel dimension and are fed into a convolutional layer. Finally, the resultant maps are split into two sets along the channel dimension. Each set is added with the original feature maps of the related stream.
3.2 IFM-Inspired Semi-Supervised Learning Framework
To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig.3), inspired by the IFM theory. This framework includes a bi-directional supervised scheme and an unsupervised scheme. The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated IFM parameters to perform UIE, and a backward-degradation module, which uses estimated parameters to degrade the reference images to their real-world degraded counterparts. As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data. Also, the generalization of the model trained is stronger than that trained using a supervised method with limited real-world data. On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect. A second set of images with different degrees of degradation are generated from these images based on the IFM. Two sets of parameters can be estimated using both sets of images respectively. Each set is utilized as the reference for the other set because their IFMs are related.
Figure 3: The proposed semi-supervised learning framework, which comprises a bi-directional supervised learning scheme and an unsupervised learning scheme, used for training our PA-UIENet.
Bi-directional Supervised Learning. Given a set of labeled training data, we first use the estimated parameters to obtain enhanced results $\hat{J}^{c}(x)$. Then we perform the forward-enhancement learning module by minimizing the following loss function:
[…]
Due to the ill-posed nature of the IFM, however, only $\mathcal{L}_{fwd}$ may be insufficient for ensuring the validity of the estimated parameters. Therefore, we propose a backward-degradation learning module which degrades clear reference images using the estimated parameters to their degraded counterparts $\hat{I}^{c}(x)$. The supervision over the backward-degradation is achieved by minimizing the loss function:
[…]
Inspired by the Retinex [11] theory, we further apply Gaussian blur to input images in order to derive the hint of the global ambient light, which is able to boost the training of the A-stream. Given the ambient light $\hat{A}^{c}$ estimated using the A-stream, we minimize the following loss function:
$\mathcal{L}_{A\text{-}sup}=||(I^{c}(x)*G-\hat{A}^{c}||_{2}^{2},$ (4)
where $G$ stands for a Gaussian kernel and * is the convolution operation. Finally, the bi-directional supervised learning scheme is conducted as:
$\mathcal{L}_{sup}=\mathcal{L}_{fwd}+\lambda_{1}\mathcal{L}_{bwd}+\lambda_{2}\mathcal{L}_{A\text{-}sup},$ (5)
where $\lambda_{1}$ and $\lambda_{2}$ are used to balance different loss functions. Compared with the supervised methods [19, 16, 17], this scheme exploits the limited labeled underwater images better.
Unsupervised Learning. To leverage unlabeled underwater images, generative networks [21, 15] were used. However, they were difficult to train and produced unstable results. Since unsupervised learning methods [23, 24] are able to directly learn from degraded images without requiring reference images, we are motivated to adopt an unsupervised learning scheme on top of the IFM.
Given an unlabeled real-world degraded image $I^{c}_{1}(x)$, the IFM can be expressed as:
$I^{c}_{1}(x)=J^{c}_{1}(x)t^{c}_{1}(x)+(1-t^{c}_{1}(x))A^{c}_{1}.$ (6)
As performed in [23], a new degraded image $I^{c}_{2}(x)$ can be purposely derived according to:
$I^{c}_{2}(x)=\alpha I^{c}_{1}(x)+(1-\alpha)A^{c}_{1},$ (7)
where $\alpha\in(0,1)$ is a controlled factor which decides the degradation extent. Substituting $I^{c}_{1}(x)$ in Eq. (6) into Eq. (7), $I^{c}_{2}(x)$ can be then expressed as:
$I^{c}_{2}(x)=J^{c}_{1}(x)(\alpha t^{c}_{1}(x))+(1-\alpha t^{c}_{1}(x))A^{c}_{1}.$ (8)
Eq. (8) also satisfies the underwater IFM but presents the more severe degradation than the $I^{c}_{1}(x)$, generated from the same underlying clear image $J^{c}_{1}(x)$, because $\alpha t^{c}_{1}(x)<t^{c}_{1}(x)$ always holds true. Therefore, the parameters $\hat{t}^{c}_{2}(x)$ and $\hat{A}^{c}_{2}$ estimated using the PA-UIENet from image $I^{c}_{2}(x)$ will be close to $\alpha\hat{t}^{c}_{1}(x)$ and $\hat{A}^{c}_{1}$ respectively, where $\hat{t}^{c}_{1}(x)$ and $\hat{A}^{c}_{1}$ are the parameters estimated from $I^{c}_{1}(x)$ [23]. Correspondingly, two loss functions are defined as:
[…]
Due to the lack of reference images, we use a non-reference loss function [24] based on the Gray-world prior [42] to constrain the enhancement of image $\hat{J}^{c}_{1}(x)$, which is expressed as:
[…]
where $E(J^{c}_{1}(x))$ denotes the average of a specific channel $c$. Finally, we define the loss function of the unsupervised learning scheme as:
$\mathcal{L}_{unsup}=\mathcal{L}_{T}+\mathcal{L}_{A\text{-}unsup}+\lambda_{3}\mathcal{L}_{gw},$ (12)
where $\lambda_{3}$ is a weighting factor.
Semi-Supervised Learning. On top of both the bi-directional supervised learning and unsupervised learning schemes, the proposed semi-supervised learning framework can be formulated as:
$\mathcal{L}_{semi\text{-}sup}=\mathcal{L}_{sup}+\lambda_{unsup}\mathcal{L}_{unsup},$ (13)
where $\lambda_{unsup}$ is used to balance the two schemes.
4 Experimental Settings
In this section, we will briefly introduce the baselines, data sets, evaluation metrics and implementation details utilized in our experiments.
Baselines. We compared the proposed method with four prior-based approaches which were developed on top of the IFM, including DCP [2], UDCP [3], GDCP [4] and DHL [8], two prior-based methods which did not consider the IFM, including Histogram Equalization (HE) and Retinex [11], and two supervised deep learning approaches, including WaterNet [16] and UColor [17].
Data Sets. We conducted a series of experiments on five testing sets derived from four publicly available real-world underwater data sets: UIEB [16], SUIM-E [18], RUIE [43] and SQUID [8]. The UIEB data set originally contains 890 pairs of degraded images, in which each reference image was selected from the results produced by applying 12 algorithms to a degraded image. Ten pairs of UIEB [16] images were discarded because the degraded images in these pairs had been included in the rest pairs. Only 880 pairs of UIEB images were used in this study. The SUIM-E data set comprises 1,525 pairs of degraded underwater images and a testing set of 110 labeled images. Three subsets are comprised of the RUIE data set, including UIQS, UCCS and UHTS, which include 3,630, 300 and 300 unlabeled images, respectively. The SQUID data set comprises 114 underwater images captured using a stereo camera, divided into the Michmoret, Katzaa, Nachsholim and Satil subsets.
Regarding the supervised learning, we randomly selected 720 pairs from the 880 pairs UIEB [16] images and used these as the training set. For the unsupervised learning, we collected 1,934 unlabeled real-world underwater images, which cover diverse underwater scenes. In total, our training set contained 2,654 images. Among the rest of the UIEB [16] data set, 80 pairs of images were randomly selected as the validation set, while the remaining pairs were used as the testing set which is referred to as Test-U80. The UIEB data set also consists of 60 challenging images without reference, which were used as the second testing set, referred to as Test-C60. A third testing set was obtained from the testing set of the SUIM-E data set [18], which contains 110 pairs of labeled underwater images and is named as Test-S110. For the RUIE [43] data set, the UCCS was utilized as the fourth testing set, denoted as Test-UCCS. Besides, 53 SQUID [8] images captured using the camera at the right hand side were randomly selected and were used as the fifth testing set, namely, Test-R53.
Evaluation Metrics. For the labeled testing data, we used both Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [44] to assess the quality of an enhanced image with regard to the reference image. Regarding the unlabeled testing data, we used the Underwater Image Quality Measure (UIQM) [45] and Underwater Color Image Quality Evaluation (UCIQE) [46] metrics. When the SQUID [8] data set was tested, the Average Reproduction Angular Error ($\bar{\psi}$) [8] was used to evaluate the quality of color restoration. Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [8].
Implementation Details. We implemented our method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3080 graphics processing unit. During the training process, the images were resized to the resolution of $256\times 256$ pixels. The number of filters in the RCT was 16. For the $Enc_{1}$ to $Dec_{1}$ in the T-stream, the numbers of filters were 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn. Regarding the A-stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6. We trained the PA-UIENet using the AdamW [47] optimizer. The learning rate and batch size were set to 1e-4 and 6 respectively. Four weighting factors, including $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{unsup}$ were set to 0.001, 0.005, 10 and 0.001 respectively. We first trained the PA-UIENet using the bi-directional supervised scheme for 3,000 iterations as a warm-up stage. Then the semi-supervised scheme was used to train it for 150,000 iterations. When the semi-supervised learning was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for every 120 iterations.
5 Experimental Results
In this section, we will report the results obtained in the underwater image enhancement evaluation, parameter estimation and color restoration evaluation and ablation study.
5.1 Evaluation on Underwater Image Enhancement
The quantitative evaluation of eight baselines and the proposed method was conducted for the UIE task. The results are reported in Tables 1. As can be seen, our method outperformed both six prior-based methods and two supervised deep learning approaches on two labeled data sets, including Test-U80 and Test-S110, when the PSNR and SSIM metrics were used. On the other hand, the UIQM and UCIQE metrics were also used in the evaluation experiment where we did not utilize reference images. In this case, the three deep learning methods were inferior to the prior-based approaches. This finding is consistent with those reported in the existing studies [16, 18]. Specifically, the performance of our method was normally comparable to that of the two deep learning approaches, i.e., WaterNet [16] and UColor [17], and was comparable to that of the prior-based methods. As demonstrated in [16, 18, 25], both the UIQM and UCIQE metrics are not always consistent with the quality of the images processed by UIE methods. It should be noted that, however, the prior-based methods did not produce stable results across the five testing sets. This should be attributed to the fact that these methods normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene is not suited to the prior assumption [4, 16].
Table 1: The quantitative evaluation of eight baselines and our method on five real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts respectively.
Method Test-U80 Test-S110 Test-C60 Test-UCCS Test-R53
[…]
DCP 12.48 63.34 1.35 0.51 14.52 67.38 1.36 0.57 0.56 0.51 0.03 0.44 -2.05 0.42
UDCP 9.37 33.62 2.69 0.59 10.04 34.28 1.23 0.60 1.24 0.54 1.83 0.57 0.76 0.59
GDCP 12.46 67.41 2.17 0.56 14.67 71.29 2.12 0.64 1.49 0.60 1.24 0.56 5.50 0.50
DHL 14.49 62.81 3.19 0.64 14.96 69.68 2.35 0.66 1.85 0.64 2.98 0.63 0.22 0.64
HE 16.48 77.77 3.58 0.67 15.37 74.76 2.97 0.68 2.32 0.65 3.93 0.67 1.71 0.64
Retinex 1038 61.25 2.33 0.48 8.29 53.44 0.85 0.45 1.21 0.46 2.03 0.47 -2.03 0.51
WaterNet 17.02 70.27 2.11 0.50 18.80 75.19 1.47 0.54 1.28 0.51 1.25 0.45 0.09 0.45
UColor 20.93 85.19 3.44 0.60 20.23 84.23 2.30 0.63 2.08 0.58 3.22 0.55 1.28 0.54
Ours 22.82 88.87 3.60 0.57 20.46 87.22 2.35 0.60 2.02 0.56 3.24 0.53 1.92 0.56
The qualitative comparison of the results produced by the eight baselines and our method is shown in Fig.4. It can be seen that the enhanced images derived using our method presented the better visual quality, compared with the images enhanced by the baselines. Our method not only effectively removed the color casts in the degraded images but also greatly increased the visibility of them. In contrast, the prior-based methods usually introduced artifacts that decreased the visual quality while the learning-based methods could not eliminate the color casts completely.
Figure 4: The results obtained using eight baselines and our method on five degraded images in Test-U80, Test-S110, Test-C60, Test-UCCS and Test-R53, respectively, shown from top to bottom.
5.2 Evaluation on Transmission Estimation and Color Restoration
The quantitative evaluation of our methods and eight baselines for transmission estimation and color restoration is reported in Table 5.2. It can be seen that our method achieved the superior, or at least the comparable, performance to that of the eight baselines on four different subsets of Test-R53. Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets no matter which metric was considered, while it performed comparably to the best baseline on the Nachsholim and Satil subsets. It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset. The inferior results may be attributed to the special scenes contained in this subset [25]. As shown in Fig. 5, our method is able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
Table 2: The quantitative evaluation of eight baselines and our method on the four subsets of Test-R53, where PCC and $\bar{\psi}$ are used for transmission estimation and color restoration respectively.
Method Michmoret Katzaa Nachsholim Satil
[…]
DCP -0.16 34.63 0.00 35.47 -0.26 34.75 0.03 36.21
UDCP -0.54 36.97 -0.15 40.43 0.06 38.79 0.12 51.45
GDCP 0.29 33.76 0.24 34.47 -0.04 34.28 -0.03 35.31
DHL -0.07 32.10 0.10 35.50 0.39 35.91 0.11 34.82
HE - 35.44 - 35.43 - 34.37 - 36.33
Retinex - 31.44 - 34.37 - 33.29 - 35.31
WaterNet - 25.40 - 25.70 - 25.80 - 28.31
UColor - 15.55 - 13.94 - 17.21 - 21.43
Ours 0.71 11.51 0.48 8.75 0.18 19.59 0.04 13.46
Table 3: The ablation study on Test-U80 and the Michmorest subset in Test-R53.
Method Test-U80 Michmoret
[…]
Simplified 22.16 89.57 -0.15
w/ RCT 22.97 89.93 0.58
w/ Bi-Sup 22.51 88.94 0.68
w/ Semi-Sup 22.82 88.87 0.71
Figure 5: Each group shows a degraded image and the red channels of five transmission maps (top) and five enhanced images (bottom) obtained using four prior-based baselines and our method.
5.3 Ablation Study
To examine the effect of each component of our method, including the RCT, bi-directional supervised scheme and semi-supervised learning framework, we conducted an ablation study. For simplicity, we only utilized Test-U80 and the Michmoret subset in Test-R53. The PA-UIENet without the RCT together with the forward-enhancement supervised learning module was crafted as a simplified version. On top this version, we derived three additional variants by inserting the RCT (w/ RCT) and then replacing the forward-enhancement supervised module by the bi-directional supervised scheme (w/ Bi-Sup) or the semi-supervised learning framework (w/ Semi-Sup). As shown in Table 5.2, the simplified version outperformed all the baselines in terms of the PSNR and SSIM metrics. However, this variant did not estimate transmission maps well, which may be attributed to the overfitting problem resulted from limited training data. When the RCT was used, the performance of image enhancement and transmission estimation was boosted. It was indicated that the adaptability of the red channel increased the robustness and generalization of the model. Though the utilization of the bi-directional-supervised learning scheme or the semi-supervised learning framework had a negative effect on the PSNR and SSIM values, the effectiveness of transmission estimation continuously increased. It is suggested that the use of the backward-degradation learning module or the unsupervised learning scheme alleviates the overfitting problem.
6 Conclusion
In this paper, we developed a novel Physics-Aware Dual-Stream Underwater Image Enhancement Network, namely, PA-UIENet, which explicitly estimates the degradation parameters of the IFM, for the UIE task. To overcome the challenge of insufficient data, we also adopted an IFM-inspired semi-supervised learning framework, comprising a bi-directional supervised scheme and an unsupervised scheme. Due to the complementary action of the two schemes, the PA-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model is stronger, compared to supervised and unsupervised methods. To our knowledge, this study made the first attempt to jointly exploit the physics-aware deep network and the IFM-inspired semi-supervised learning technique for the UIE task. The proposed method performed better than, or at least comparably to, eight baselines on five real-world underwater testing sets in different evaluations. The promising results should be due to the fact that our method is able to not only estimate degradation parameters but also learn the characteristics of diverse underwater scenes.